# BatteryLife: A Comprehensive Dataset and Benchmark for Battery Life Prediction

### v1

@@ Abstract.
Battery Life Prediction (BLP), which relies on time series data produced by battery degradation tests, is crucial for battery utilization, optimization, and production. Despite impressive advancements, this research area faces three key challenges. Firstly, the limited size of existing datasets impedes insights into modern battery life data. Secondly, most datasets are restricted to small-capacity lithium-ion batteries tested under a narrow range of diversity in labs, raising concerns about the generalizability of findings. Thirdly, inconsistent and limited benchmarks across studies obscure the effectiveness of baselines and leave it unclear if models popular in other time series fields are effective for BLP. To address these challenges, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. BatteryLife integrates 16 datasets, offering a 2.4 times sample size compared to the previous largest dataset, and provides the most diverse battery life resource with batteries from 8 formats, 80 chemical systems, 12 operating temperatures, and 646 charge/discharge protocols, including both laboratory and industrial tests. Notably, BatteryLife is the first to release battery life datasets of zinc-ion batteries, sodium-ion batteries, and industry-tested large-capacity lithium-ion batteries. With the comprehensive dataset, we revisit the effectiveness of baselines popular in this and other time series fields. Furthermore, we propose CyclePatch, a plug-in technique that can be employed in a series of neural networks. Extensive benchmarking of 18 methods reveals that models popular in other time series fields can be unsuitable for BLP, and CyclePatch consistently improves model performance establishing state-of-the-art benchmarks. Moreover, BatteryLife evaluates model performance across aging conditions and domains. BatteryLife is available at https://github.com/Ruifeng-Tan/BatteryLife.
time series, battery, AI for chemistry, materials informatics
@@ 1. INTRODUCTION
Rechargeable batteries are ubiquitous in modern industry, including electric vehicles, power grids, and portable devices (Tarascon and Armand, 2010; Huang et al., 2022; Tao et al., 2023a; Lu et al., 2022). Nevertheless, batteries inevitably degrade with cyclic operation due to intrinsic electrochemical mechanisms (Birkl et al., 2017; Downie et al., 2014; Tan et al., 2024). The degradation causes shorter usable time and may lead to safety issues in the applications (Zhang et al., 2024; Zhao et al., 2024; Zhang et al., 2023). Typically, batteries are deemed at end-of-life when their capacity falls to 80% of the initial, and end-of-life batteries are unsuitable for high-demanding applications like electric vehicles (Severson et al., 2019; Attia et al., 2020; Ma et al., 2022). To ensure more sustainable and safer battery operation, battery engineers/scientists need to conduct degradation tests to measure the battery life during battery optimization and quality control (Attia et al., 2020; Li et al., 2024; Stock et al., 2022). However, degradation tests are time-consuming since non-linear capacity loss occurs over months to years before the end of life. Accurate early prediction of the battery life is thus vital for battery utilization, optimization, and production.
Machine learning models have been developed as promising approaches for Battery Life Prediction (BLP) (Severson et al., 2019; Hu et al., 2020; Jiang et al., 2021; Weng et al., 2021; Ma et al., 2022; Zhou and Zhang, 2024; Hsu et al., 2022; Zhang et al., 2018; Fei et al., 2023; Zhang et al., 2025a; Ren et al., 2021; Song et al., 2018; Tao et al., 2023b). In early works, battery experts generally made arduous efforts to extract features from the voltage as well as current time series of degradation tests (Severson et al., 2019; Jiang et al., 2021) and data of additional tests such as temperature test (Zhang et al., 2025b) and acoustic test (Knehr et al., 2018) based on domain knowledge. The extracted features are then fed into classic machine learning models such as Gaussian process regressor to predict the battery life. Recently, with the advent of larger datasets, neural networks have shown advantages over classic machine learning models without requiring deep domain expertise (Ma et al., 2022; Xu et al., 2021; Tang et al., 2022; Fei et al., 2023; Zhang et al., 2025a; Tao et al., 2023b). Despite the success these methods have achieved, they also reveal three issues that impede advances in this field:
@@
Dataset Battery/Sample Number Format Chemical System Operation Temperature Protocol Data Source Battery Type Year
CALCE (Xing et al., 2013) 13/1,300 1 1 1 2 Lab test Li-ion 2013
HNEI (Devie et al., 2018) 14/1,400 1 1 1 2 Lab test Li-ion 2018
MATR (Attia et al., 2020; Severson et al., 2019) 169/16,900 1 1 1 81 Lab test Li-ion 2019,2020
UL_PUR (Juarez-Robles et al., 2020, 2021) 10/1,000 1 1 1 1 Lab test Li-ion 2020
SNL (Preger et al., 2020) 52/5,200 1 3 3 23 Lab test Li-ion 2020
MICH_EXP (Mohtat et al., 2021) 18/1,800 1 1 3 6 Lab test Li-ion 2021
MICH (Weng et al., 2021) 40/4,000 1 1 2 2 Lab test Li-ion 2021
RWTH (Li et al., 2021) 48/4,800 1 1 1 1 Lab test Li-ion 2021
HUST (Ma et al., 2022) 77/7,700 1 1 1 77 Lab test Li-ion 2022
Tongji (Zhu et al., 2022) 108/10,800 1 1 3 6 Lab test Li-ion 2022
BatteryArchive (bat, 2024) 147/14,700 3 5 6 36 Lab test Li-ion 2024
Stanford (Cui et al., 2024) 41/4,100 1 1 3 1 Lab test Li-ion 2024
XJTU (Wang et al., 2024c) 23/2,300 1 1 1 2 Lab test Li-ion 2024
ISU_ILCC (Li et al., 2024) 240/24,000 1 1 1 202 Lab test Li-ion 2024
BatteryML (Zhang et al., 2024) 403/40,300 2 5 5 189 Lab test Li-ion 2024
BatteryLife 998/99,800 8 80 12 646 Lab test; Industrial test Li-ion; Zn-ion; Na-ion 2025
  • Limited dataset size. As shown in Table 1, the previous largest dataset is BatteryML (Zhang et al., 2024) that offers 403 end-of-life batteries by integrating 7 public datasets (Attia et al., 2020; Severson et al., 2019; Ma et al., 2022; Devie et al., 2018; Preger et al., 2020; Juarez-Robles et al., 2020; Li et al., 2021; Juarez-Robles et al., 2021). However, this undertaking overlooks several datasets (Zhu et al., 2022; Weng et al., 2021; Mohtat et al., 2021) and lacks datasets (Cui et al., 2024; Wang et al., 2024c; Li et al., 2024) that are released recently. In this vein, BatteryML, whose size is smaller than half of the sum of existing datasets, fails to provide comprehensive insights into modern battery life data.
  • Limited data diversity. According to Table 1, most existing battery life datasets offer only lithium-ion (Li-ion) batteries with one format, one chemical system, one operation temperature, and several charge/discharge protocols. The BatteryML (Zhang et al., 2024) and ISU_ILCC (Li et al., 2024) offer great diversity in charge/discharge protocols, but their diversity in other aspects is still limited. Researchers (Ma et al., 2022) have observed that models validated on various charge protocols may be ineffective in learning varying discharge protocols. This raises concerns about the generalizability of findings derived from datasets with limited diversity. Furthermore, the limited diversity results in a lack of data needed to inform other battery research directions.
@@
To address these issues, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. From a dataset perspective, BatteryLife stands out as the largest and most diverse battery life dataset by integrating 16 datasets. Specifically, BatteryLife is 2.4 times the size of BatteryML (Zhang et al., 2024) in terms of battery number, covering more than 90,000 samples. Remarkably, BatteryLife furnishes unparalleled diversity, delivering 4 times format, 16 times chemical system, 2.4 times operating temperature, and 3.4 times charge/discharge protocol compared to BatteryML (Zhang et al., 2024). Notably, BatteryLife pioneers as the first work that releases zinc-ion (Zn-ion) and sodium-ion (Na-ion) battery life datasets, and additionally offers large-capacity Li-ion batteries that are tested by the battery manufacturer (i.e., CALB Group Co., Ltd.). The formidable diversity challenges the model’s capability to learn essential mappings from early data to life and allows inspiration for the data mining community and a broader battery community.
@@
With respect to benchmark, BatteryLife provides a unified experimental setup and supports BLP across cycles ranging from one to one hundred, covering a wide range of application scenarios. Moreover, BatteryLife serves as a comprehensive benchmark for BLP. It encompasses methods popular in this and other time series fields. Additionally, motivated by the patching technique (Nie et al., 2023) and the inherent characteristics of degradation tests, we propose CyclePatch, a plug-in technique for modeling degradation test data. As illustrated in Figure 1, voltage and current time series exhibit similar patterns across cycles within a protocol. This observation motivates us to treat each cycle as a token in CyclePatch, explicitly capturing recurring patterns in degradation tests. Furthermore, Figure 1 demonstrates that cycling data from batteries with different life labels show discriminative characteristics, both in terms of cycles of the same number and variance across cycles. Thus, CyclePatch employs an intra-cycle encoder to model the complex interactions among variables within each cycle. In this framework, CyclePatch generates informed representations for each cycle, on which an inter-cycle encoder is applied to learn effective patterns across cycle tokens. Through extensive experiments of 18 methods, we reveal that some techniques widely adopted in other time series fields are inappropriate for BLP. Additionally, models consistently achieve improved performance when CyclePatch is integrated. Moreover, the intricacies of achieving effective BLP are further disclosed via evaluations across aging conditions and domains.
@@
  • The largest battery life dataset: BatteryLife offers more than 90,000 samples from 998 batteries, which is 2.4 times the size of the previous largest one. The unprecedented dataset size allows researchers to generate practical insights into battery life data by making advances on BatteryLife.
@@ 2.1. Degradation Test
where $Q_{i}$ is the capacity of $i^{th}$ cycle, and $Q_{0}$ is the initial capacity. The literature (Severson et al., 2019; Ma et al., 2022; Zhang et al., 2024; Wang et al., 2024c; Roman et al., 2021; Lu et al., 2023) typically takes discharge capacity as $Q_{i}$. Regarding $Q_{0}$, some works (Severson et al., 2019; Ma et al., 2022; Zhang et al., 2024; Wang et al., 2024c) utilize nominal capacity as $Q_{0}$ while some (Roman et al., 2021; Lu et al., 2023) set $Q_{1}$ as $Q_{0}$. Non-linear SOH degradation can occur hundreds of cycles before batteries reach the end of life (80% SOH) as shown in Figure 3.
@@ 3.1. Dataset Overview
BatteryLife dataset is constructed via integration of 3 our datasets and 13 previously available datasets (Xing et al., 2013; Devie et al., 2018; Attia et al., 2020; Severson et al., 2019; Juarez-Robles et al., 2020; Preger et al., 2020; Li et al., 2021; Weng et al., 2021; Mohtat et al., 2021; Ma et al., 2022; Zhu et al., 2022; Cui et al., 2024; Wang et al., 2024c; Li et al., 2024). The data statistics of BatteryLife and previous datasets are concluded in Table 1. According to Table 1, the previous public datasets contain only Li-ion batteries collected from laboratory tests. We supply Zn-ion and Na-ion batteries by releasing batteries tested in our lab and provide industry-tested large-capacity Li-ion batteries accessed from CALB Group Co., Ltd.. On this note, BatteryLife features as the largest and most diverse battery life dataset. Note that statistics for all datasets are conducted after data preprocessing. We refer readers to Appendix A for the details of data preprocessing, simple introductions to each dataset, and degradation test details of 3 our datasets. For convenient usage, we standardize all datasets in a pickle file format with unified naming rules following BatteryML (Zhang et al., 2024), whereas the original data come from various naming rules and file formats.
@@
Figure 4. Life label distribution of BatteryLife.
@@ 3.2. Popular Benchmark Methods
We set up a series of benchmark methods in BatteryLife. First, we select methods that are widely adopted for BLP (Fei et al., 2023; Hsu et al., 2022; Ma et al., 2022; Zhang et al., 2025a; Ren et al., 2021; Song et al., 2018), including Transformer encoder (Vaswani et al., 2017), Long short-term memory (LSTM), bidirectional long short-term memory (BiLSTM), gate recurrent unit (GRU), bidirectional gate recurrent unit (BiGRU), convolutional neural network (CNN), and multilayer perceptron (MLP). Among these models, Transformer encoder, LSTM, BiLSTM, GRU, and BiGRU are directly applied to process input that has a variable length, and then a linear projection is used to produce the prediction. For CNN, we pad the input to the same length using zeros and then reshape the input into an image-like matrix using voltage, current, and capacity as three channels. Specifically, the reshaped input has a dimension of (3, 100, 300), where 100 is the largest input cycle number, and 300 is the number of resampled points within a cycle. With the reshaped input, 2-dimension CNN can be used to model cycling data and a linear projection is employed to produce predictions based on the flattened output of the CNN. As for MLP, we utilize it to model the flattened zero-padded cycling data with a dimension of ($3\times T$) and set its last layer as the output layer.
Additionally, considering cycling data are multivariate time series, we also study the effectiveness of models (Wang et al., 2023; Nie et al., 2023; Wu et al., 2021; Zeng et al., 2023; Liu et al., 2024), which are originally developed for other multivariate time series, for BLP. Specifically, we further include five baselines: DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2023), Autoformer (Wu et al., 2021), iTransformer (Liu et al., 2024) and MICN (Wang et al., 2023). The implementation details for the popular benchmark methods are described in Appendix B.2.

### v3

@@ Abstract.
Battery Life Prediction (BLP), which relies on time series data produced by battery degradation tests, is crucial for battery utilization, optimization, and production. Despite impressive advancements, this research area faces three key challenges. Firstly, the limited size of existing datasets impedes insights into modern battery life data. Secondly, most datasets are restricted to small-capacity lithium-ion batteries tested under a narrow range of diversity in labs, raising concerns about the generalizability of findings. Thirdly, inconsistent and limited benchmarks across studies obscure the effectiveness of baselines and leave it unclear if models popular in other time series fields are effective for BLP. To address these challenges, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. BatteryLife integrates 16 datasets, offering a 2.5 times sample size compared to the previous largest dataset, and provides the most diverse battery life resource with batteries from 8 formats, 59 chemical systems, 9 operating temperatures, and 421 charge/discharge protocols, including both laboratory and industrial tests. Notably, BatteryLife is the first to release battery life datasets of zinc-ion batteries, sodium-ion batteries, and industry-tested large-capacity lithium-ion batteries. With the comprehensive dataset, we revisit the effectiveness of baselines popular in this and other time series fields. Furthermore, we propose CyclePatch, a plug-in technique that can be employed in various neural networks. Extensive benchmarking of 18 methods reveals that models popular in other time series fields can be unsuitable for BLP, and CyclePatch consistently improves model performance establishing state-of-the-art benchmarks. Moreover, BatteryLife evaluates model performance across aging conditions and domains. BatteryLife is available at https://github.com/Ruifeng-Tan/BatteryLife.
time series; battery; AI for chemistry; materials informatics
@@ 1. INTRODUCTION
Rechargeable batteries are ubiquitous in modern industry, including electric vehicles, power grids, and portable devices (Huang et al., 2022; Tao et al., 2023a; Lu et al., 2022). Nevertheless, batteries inevitably degrade with cyclic operation due to intrinsic electrochemical mechanisms (Birkl et al., 2017; Downie et al., 2014; Tan et al., 2024). The degradation causes shorter usable time and may lead to safety issues in the applications (Zhang et al., 2023a, b). Typically, batteries are deemed at end-of-life when their capacity falls to 80% of the initial, and end-of-life batteries are unsuitable for high-demanding applications like electric vehicles (Severson et al., 2019; Attia et al., 2020; Ma et al., 2022). To ensure more sustainable and safer battery operation, battery engineers/scientists need to conduct degradation tests to measure the battery life during battery optimization and quality control (Attia et al., 2020; Li et al., 2024; Stock et al., 2022). However, degradation tests are time-consuming since non-linear capacity loss occurs over months to years before the end of life. Accurate early prediction of the battery life is thus vital for battery utilization, optimization, and production.
Machine learning models have been developed as promising approaches for Battery Life Prediction (BLP) (Severson et al., 2019; Hu et al., 2020; Jiang et al., 2021; Weng et al., 2021; Ma et al., 2022; Zhou and Zhang, 2024; Hsu et al., 2022; Zhang et al., 2018; Fei et al., 2023; Zhang et al., 2025a; Ren et al., 2021; Tao et al., 2023b). In early works, battery experts generally made arduous efforts to extract features from the voltage as well as current time series of degradation tests (Severson et al., 2019; Jiang et al., 2021) and data of additional tests such as temperature test (Zhang et al., 2025b) and acoustic test (Knehr et al., 2018) based on domain knowledge. The extracted features are then fed into classic machine learning models such as Gaussian process regressor to predict the battery life. Recently, with the advent of larger datasets, neural networks have shown advantages over classic machine learning models without requiring deep domain expertise (Ma et al., 2022; Xu et al., 2021; Tang et al., 2022; Fei et al., 2023; Zhang et al., 2025a; Tao et al., 2023b). Despite the success these methods have achieved, they also reveal three issues that impede advances in this field:
@@
Dataset Battery Number Format Chemical System Operation Temperature Protocol Data Source Battery Type Year
CALCE (Xing et al., 2013) 13 1 1 1 2 Lab test Li-ion 2013
HNEI (Devie et al., 2018) 14 1 1 1 2 Lab test Li-ion 2018
MATR (Attia et al., 2020; Severson et al., 2019) 169 1 1 1 81 Lab test Li-ion 2019,2020
UL_PUR (Juarez-Robles et al., 2020, 2021) 10 1 1 1 1 Lab test Li-ion 2020
SNL (Preger et al., 2020) 50 1 3 3 23 Lab test Li-ion 2020
MICH_EXP (Mohtat et al., 2021) 18 1 1 3 6 Lab test Li-ion 2021
MICH (Weng et al., 2021) 40 1 1 2 2 Lab test Li-ion 2021
RWTH (Li et al., 2021) 48 1 1 1 1 Lab test Li-ion 2021
HUST (Ma et al., 2022) 77 1 1 1 77 Lab test Li-ion 2022
Tongji (Zhu et al., 2022) 108 1 3 3 6 Lab test Li-ion 2022
BatteryArchive (bat, 2024) 139 3 8 6 36 Lab test Li-ion 2024
Stanford (Cui et al., 2024) 41 1 1 1 1 Lab test Li-ion 2024
XJTU (Wang et al., 2024c) 23 1 1 1 2 Lab test Li-ion 2024
ISU_ILCC (Li et al., 2024) 240 1 1 1 202 Lab test Li-ion 2024
BatteryML (Zhang et al., 2023a) 381 2 5 5 186 Lab test Li-ion 2024
BatteryLife 990 8 59 9 421 Lab test; Industrial test Li-ion; Zn-ion; Na-ion 2025
  • Limited dataset size. As shown in Table 1, the previous largest dataset is BatteryML (Zhang et al., 2023a) that offers 381 end-of-life batteries by integrating 7 public datasets (Attia et al., 2020; Severson et al., 2019; Ma et al., 2022; Devie et al., 2018; Preger et al., 2020; Juarez-Robles et al., 2020; Li et al., 2021; Juarez-Robles et al., 2021). However, this undertaking overlooks several datasets (Zhu et al., 2022; Weng et al., 2021; Mohtat et al., 2021) and lacks datasets (Cui et al., 2024; Wang et al., 2024c; Li et al., 2024) that are released recently. In this vein, BatteryML, whose size is smaller than half of the sum of existing datasets, fails to provide comprehensive insights into modern battery life data.
  • Limited data diversity. According to Table 1, most existing battery life datasets offer only lithium-ion (Li-ion) batteries with one format, one chemical system, one operation temperature, and several charge/discharge protocols. The BatteryML (Zhang et al., 2023a) and ISU_ILCC (Li et al., 2024) offer great diversity in charge/discharge protocols, but their diversity in other aspects is still limited. Researchers (Ma et al., 2022) have observed that models validated on various charge protocols may be ineffective in learning varying discharge protocols. This raises concerns about the generalizability of findings derived from datasets with limited diversity. Furthermore, the limited diversity results in a lack of data needed to inform broader battery research directions.
@@
To address these issues, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. From a dataset perspective, BatteryLife stands out as the largest and most diverse battery life dataset by integrating 16 datasets. Specifically, BatteryLife is 2.5 times the size of BatteryML (Zhang et al., 2023a) in terms of battery number, covering more than 90,000 samples. Remarkably, BatteryLife furnishes unparalleled diversity, delivering 4 times format, 11.8 times chemical system, 1.8 times operation temperature, and 2.2 times charge/discharge protocol compared to BatteryML (Zhang et al., 2023a). Notably, BatteryLife pioneers as the first work that releases zinc-ion (Zn-ion) and sodium-ion (Na-ion) battery life datasets, and additionally offers large-capacity Li-ion batteries that are tested by the battery manufacturer (i.e., CALB Group Co., Ltd.). The formidable diversity challenges the model’s capability to learn essential mappings from early data to life and allows inspiration for the data mining community and a broader battery community.
@@
[…]
[zero figure]Partial cycling data (voltage and current time series) from three batteries with the same specifications tested under a protocol.
With respect to benchmark, BatteryLife provides a unified experimental setup and supports BLP across cycles ranging from one to one hundred, covering a wide range of application scenarios. Moreover, BatteryLife serves as a comprehensive benchmark for BLP. It encompasses methods popular in this and other time series fields. Additionally, motivated by the patching technique (Nie et al., 2023) and the inherent characteristics of degradation tests, we propose CyclePatch, a plug-in technique for modeling degradation test data. As illustrated in Figure 1, for batteries with the same specifications, voltage and current time series exhibit similar patterns across cycles within a protocol. This observation motivates us to treat each cycle as a token in CyclePatch, explicitly capturing recurring patterns in degradation tests. Furthermore, Figure 1 demonstrates that cycling data from batteries with different life labels show discriminative characteristics, both in terms of cycles of the same number and variance across cycles. Thus, CyclePatch employs an intra-cycle encoder to model the complex interactions among variables within each cycle. In this framework, CyclePatch generates informed representations for each cycle, on which an inter-cycle encoder is applied to learn effective patterns across cycle tokens. Through extensive experiments of 18 methods, we reveal that some techniques widely adopted in other time series fields are inappropriate for BLP. Additionally, models consistently achieve improved performance when CyclePatch is integrated. Moreover, the intricacies of achieving effective BLP are further disclosed via evaluations across aging conditions and domains.
@@
  • The largest battery life dataset: BatteryLife offers more than 90,000 samples from 990 batteries, which is 2.5 times the size of the previous largest one. The unprecedented dataset size allows researchers to generate practical insights into battery life data by making advances on BatteryLife.
@@ 2.1. Degradation Test
where $Q_{i}$ is the capacity of $i^{th}$ cycle, and $Q_{0}$ is the initial capacity. The literature (Severson et al., 2019; Ma et al., 2022; Zhang et al., 2023a; Wang et al., 2024c; Roman et al., 2021; Lu et al., 2023) typically takes discharge capacity as $Q_{i}$. Regarding $Q_{0}$, some works (Severson et al., 2019; Ma et al., 2022; Zhang et al., 2023a; Wang et al., 2024c) utilize nominal capacity as $Q_{0}$ while some (Roman et al., 2021; Lu et al., 2023) set $Q_{1}$ as $Q_{0}$. Non-linear SOH degradation can occur hundreds of cycles before batteries reach the end of life (80% SOH) as shown in Figure 3.
@@
[…]
[first figure]Partial cycling data in degradation tests from (a) a lab-tested Li-ion battery, (b) a lab-tested Zn-ion battery, (c) a lab-tested Na-ion battery, and (d) an industry-tested Li-ion battery. The positive and negative values in current represent charging and discharging, respectively. The gray dotted lines are separators of different cycles. The charging and discharging periods are labeled in the first cycle as illustrations.
@@ 3.1. Dataset Overview
BatteryLife dataset is constructed via integration of 3 our datasets and 13 previously available datasets (Xing et al., 2013; Devie et al., 2018; Attia et al., 2020; Severson et al., 2019; Juarez-Robles et al., 2020; Preger et al., 2020; Li et al., 2021; Weng et al., 2021; Mohtat et al., 2021; Ma et al., 2022; Zhu et al., 2022; Cui et al., 2024; Wang et al., 2024c; Li et al., 2024). The data statistics of BatteryLife and previous datasets are concluded in Table 1. According to Table 1, the previous public datasets contain only Li-ion batteries collected from laboratory tests. We supply Zn-ion and Na-ion batteries by releasing batteries tested in our lab and provide industry-tested large-capacity Li-ion batteries accessed from CALB Group Co., Ltd.. On this note, BatteryLife features as the largest and most diverse battery life dataset. Note that statistics for all datasets are conducted after data preprocessing. We refer readers to Appendix A for the details of data preprocessing, simple introductions to each dataset, and degradation test details of 3 our datasets. For convenient usage, we standardize all datasets in a pickle file format with unified naming rules following BatteryML (Zhang et al., 2023a), whereas the original data come from various naming rules and file formats.
@@
[…]
[SOH trajectories]SOH degradation trajectories in BatteryLife.
@@ 3.2. Popular Benchmark Methods
We set up a series of benchmark methods in BatteryLife. First, we select methods that are widely adopted for BLP (Fei et al., 2023; Hsu et al., 2022; Ma et al., 2022; Zhang et al., 2025a; Ren et al., 2021; Liu et al., 2021), including Transformer encoder (Vaswani et al., 2017), Long short-term memory (LSTM), bidirectional long short-term memory (BiLSTM), gate recurrent unit (GRU), bidirectional gate recurrent unit (BiGRU), convolutional neural network (CNN), and multilayer perceptron (MLP). Among these models, Transformer encoder, LSTM, BiLSTM, GRU, and BiGRU are directly applied to process input that has a variable length, and then a linear projection is used to produce the prediction. For CNN, we pad the input to the same length using zeros and then reshape the input into an image-like matrix using voltage, current, and capacity as three channels. Specifically, the reshaped input has a dimension of (3, 100, 300), where 100 is the largest input cycle number, and 300 is the number of resampled points within a cycle. With the reshaped input, 2-dimension CNN can be used to model cycling data and a linear projection is employed to produce predictions based on the flattened output of the CNN. As for MLP, we utilize it to model the flattened zero-padded cycling data with a dimension of ($3\times T$) and set its last layer as the output layer.
Additionally, considering cycling data are multivariate time series, we also study the effectiveness of models (Wang et al., 2023; Nie et al., 2023; Wu et al., 2021; Zeng et al., 2023; Liu et al., 2024), which are originally developed for other multivariate time series, for BLP. Specifically, we further include five baselines: DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2023), Autoformer (Wu et al., 2021), iTransformer (Liu et al., 2024) and MICN (Wang et al., 2023). The implementation details for the popular benchmark methods are described in Appendix B.2. A dummy model that uses the mean of training labels as the prediction is also included as a baseline.